INFRARED SPECTRA OF ^{12}C^{16}O IN ABSORPTION AND EVALUATION OF RADIAL FUNCTIONS FOR POTENTIAL ENERGY AND ELECTRIC DIPOLAR MOMENT

\underline{J. F. OGILVIE}, Centre for Experimental and Constructive Mathematics, Department of Mathematics, Simon Fraser University, 8888 University drive, Burnaby, British Columbia V5A 1S6 Canada; SWEE-LAN CHEAH, YUAN-PERN LEE, Department of Chemistry, National Tsing Hua University, Hsinchu 30013, Taiwan; S. P. A. SAUER, Department of Chemistry, University of Copenhagen, Universitetsparken 5, DK 2100 Copenhagen O, Denmark.